OpenStudy (anonymous):

Solve for X 8x - 2=6x + 4

OpenStudy (anonymous):

8x-6x=4+2 regroup and simplify

OpenStudy (anonymous):

8x-2=6x+4 8x-6x=4+2 2x=6 x=6/2 x=3